# A parametric study of the pipeline hammer phenomenon in plastic Bingham slurry flows using the finite element method

### v1

@@ 3.4 Effects of the gradual valve closure
An additional brief test is performed to assess the impact of the gate closure model on the wave dynamics. We set the Dirichlet boundary condition in (2.14) as
$g^{*}(t)=\frac{16}{f\widetilde{\text{Re}}}\left(1-\frac{t^{*}}{T_{c}}\right),$
where we check several closure times scaling it to the characteristic time $L/c$. The results, using the parametric configuration for the copper slurry, are shown in Figure 11 for velocity, and Figure 12, for the pressure. We depict four scenarios, namely an instant closure time ($T_{c}=0$), a fast closure time ($T_{c}=0.5$), a critical closure time ($T_{c}=1$) and a slow closure time ($T_{c}=1.5$). As expected, the sharpest numerical solutions of the governing equations are retrieved for the instantaneous closure, and we observe how the delay of the valve operation softens the overall dynamics, still keeping nonetheless similar over-pressure values. Yet, not-sudden pressure changes, as with the case $T_{c}=1.5$, are preferable operation conditions.
@@
The valve closure time study is summarized by looking at the plots of Figure 13, where we see how the overall simulation pressure peak (i.e. the first peak) behaves against both, the Hedstrom number, and the adimensional closure time. We first confirm that the Newtonian case brings higher pressure peaks, and we also see that the slow closure have a bigger consequence at the pipe half length control point. That is to say, at the valve closure control point, the peak pressure behavior is almost independent of the closure time, whereas at the middle control point, we observe a clear separation at the critical time $T_{c}=1$, after which the pressure peak decreases almost linearly.
@@ 4 Conclusions
This study provides a comprehensive analysis of the pipeline hammer phenomenon in non-Newtonian Bingham plastic fluids, employing an efficient computationally inexpensive finite element method (FEM) to accurately simulate the transient dynamics. By integrating factors such as yield stress, convective effects, and varying valve closure times, this work advances the understanding of how non-Newtonian rheology affects pressure shockwaves in pipeline systems. The findings underscore some insights:
@@
  • Valve Closure Dynamics: The analysis of valve closure time revealed that gradual valve closures can lead to delayed, slightly slower, and smooth pressure wave and pressure peaks. The detailed parametric analysis summarized in Figure 13 shows how the peak pressure, a critical design quantity, behaves for a battery of configurations.
@@
  • Adaptive Model for Laminar and Turbulent Flows: The adaptive model used for both laminar and turbulent flow regimes demonstrated that hybrid approaches incorporating finite element schemes allow for accurate predictions across various flow conditions. This adaptability is essential when dealing with fluids that exhibit both Newtonian and non-Newtonian behaviors depending on flow rates and operational conditions, which contrast to regular studies [24] where laminar friction models are adopted even in high velocity conditions, as those usually reached in transient phenomena.
@@
Overall, this work highlights the Bingham slurry parametric study, and also the advantages of using advanced numerical techniques, such as FEM, to accurately capture the complex interplay of rheological properties, flow dynamics, and operational changes in pipelines transporting non-Newtonian fluids. Future research could further refine these models through machine learning approaches for rapid parameter optimization, and stabilization techniques could be explored to enhance model robustness. This work lays a foundation for more resilient pipeline designs and informed operational strategies, particularly in industries where non-Newtonian fluids pose unique challenges to system stability and durability.

### v2

@@ 3.4 Effects of the gradual valve closure
An additional brief test is performed to assess the impact of the gate closure model on the wave dynamics. We set the Dirichlet boundary condition in (2.15) as:
$g^{*}(t^{*})=\left(1-\frac{t^{*}}{T^{*}_{c}}\right),$ (3.1)
where we check several closure times scaling it to the wave round-trip time $2L/c$. We consider the dimensionless closure time $T^{*}_{c}$, meaning that $T^{*}_{c}<2$ will lead to a fast valve closure. The results, using the parametric configuration for the copper slurry, are shown in Figure 11 for velocity, and Figure 12, for the pressure. We depict four scenarios, namely an instant closure time ($T^{*}_{c}=0$), a fast closure time ($T^{*}_{c}=1$), a critical closure time ($T^{*}_{c}=2$) and a slow closure time ($T^{*}_{c}=3$). As expected, the sharpest numerical solutions of the governing equations are retrieved for the instantaneous closure, and we observe how the delay of the valve operation softens the overall dynamics, still keeping nonetheless similar over-pressure values in the first peak, yet decreasing at the sub-sequent roundtrips. Yet, not-sudden pressure changes, as with the case $T^{*}_{c}=3$, are preferable operation conditions.
@@
The fast valve closure time study is summarized by looking at the plots of Figure 13, where we see how the overall simulation pressure peak (i.e. the first peak) behaves against both, the Hedstrom number, and the dimensionless closure time. We first confirm that the Newtonian case brings higher pressure peaks, and we also see that the critical closure reduces considerably the wave peak at the pipe half length control point ($x^{*}=1/2$). That is to say, at the valve closure control point, the peak pressure behavior is almost independent of the closure time in the fast closure setup, whereas at the middle control point, we observe a clear separation at the critical time $T_{c}=1$, after which the pressure peak decreases almost linearly.
3.5 Sensitivity to non-dimensional scaling and convective effects
We perform an additional study without assuming a specific application set-up. For this experiment, we employ $\widetilde{\text{Re}}=5000$ and $\text{He}=10^{5}$. In Figure 14 we observe the behavior of the dimensionless solutions with respect to the parameter $\lambda_{2}$ in Eqs. (2.15), keeping $\lambda_{1}=(2/10)\lambda_{2}$ (thus fixing the Mach number). One should notice that $\lambda_{2}=\delta/32\widetilde{\text{Re}}\widetilde{Ma}^{-1}=c\rho_{m}U_{0}/P_{0}$, that is to say, it gives us an idea of how much the pipe hammer over-pressure (calculated with the surrogate Joukowsky formula using the steady Newtonian velocity) exceeds the tank pressure $P_{0}$.
Figure 14 shows how $\lambda_{2}=1$ leads to an over-pressure which is not significantly different from the operational pressure induced by the tank, yielding a non-oscillatory regime, where the hammer wave is likely irrelevant. On the contrary, the case $\lambda_{2}=10$ leads to a higher pressure rise, in which the frictional forces are not able to dissipate the wave as quickly as with lower values. Any larger $\lambda_{2}$ will lead to slower dissipation rates such that friction losses become negligible (as it would be, e.g., for $\lambda_{2}=100$). This is an expected behavior, in spite of the additional resistance imposed by the plastic behavior of the fluid.
(a) Dimensionless velocity evolution over dimensionless time
(b) Dimensionless pressure evolution over dimensionless time
Figure 14: Temporal evolution of the dimensionless solutions of the transient. We move the parameter $\lambda_{1}$ from 1, where the pressure rise is comparable with the pipe operational conditions (leading to a failed pipe hammer), to 10, where the first pressure peak
On a different avenue, we fix $\lambda_{2}=10$ and compare simulations for different $\lambda_{1}$, therefore changing the relative contribution of the convective term $u^{*}\frac{\partial u^{*}}{\partial x^{*}}$ in Eqs. (2.15), and thereby also changing the coefficient $\lambda_{2}/\lambda_{1}=Ma$. We acknowledge that, even though we keep a subsonic regime, our model is unable to reproduce the fully nonlinear equations required by such conditions. This is a step further in our research, where the convective effects are indeed relevant, as shown even in Fig. 15 for our surrogate model. These differences between considering or not the convective transport effects, reveal not only a change in the peak pressure of the solutions, but also a temporal shift in their evolution, suggesting larger dissipation times, even at moderately low Mach numbers, as with the cases $\lambda_{1}=2$ or $\lambda_{1}=4$. We present this as a step towards better models for highly non-linear regimes, yet admittedly simplistic for highly compressible scenarios. We emphasize that the methodological proposal of our work (i.e. using FEM) allows this without significant changes in the numerical formulation. This would not be possible with classical methods such as the MOC.
Figure 15: Assessment of the solutions at control points with and without neglecting the convective effects in (2.15). We observe how larger differences arise as the ratio $\lambda_{1}/\lambda_{2}$ increases.
@@ 4 Conclusions
This study provides a parametric analysis of the pipeline hammer phenomenon in non-Newtonian Bingham plastic fluids, employing an efficient computationally inexpensive finite element method (FEM) to accurately simulate the transient dynamics. By integrating the effects of different physical parameters such as yield stress, convective effects, and varying valve closure times, this work advances the understanding of how non-Newtonian rheology affects pressure shockwaves in pipeline systems. The findings underscore some insights:
@@
  • Valve Closure Dynamics: The analysis of valve closure time revealed that gradual valve closures can lead to delayed, slightly attenuated, and smoother pressure waves and peaks. The parametric analysis is summarized in Figures 11 and 12 for instantaneous (ideal), fast, and slow closures. Additionally, we have investigated the variation of peak pressure across configurations in Figure 13, posing our interest in the fast closure scenario for several Bingham fluids.
@@
  • Adaptive Model for Laminar and Turbulent Flows: The adaptive quasi-steady model used for both laminar and turbulent flow regimes demonstrated that hybrid approaches can be easily incorporated in a finite element formulation. This adaptability is essential when dealing with fluids that exhibit both Newtonian and non-Newtonian behaviors depending on flow rates and operational conditions, which contrast to regular studies [26] where laminar friction models are adopted even in high velocity conditions, as those usually reached in transient phenomena.
  • Influence of Convective Effects and Dimensionless Groups: The dimensionless analysis in Section 3.5 shows that increasing the parameter $\lambda_{1}$ enhances the relative importance of convective transport. This leads to higher peak pressures and a noticeable temporal shift in the pressure response. Even in subsonic regimes, these effects are non-negligible, emphasizing the importance of including convective terms when modeling fast transients. This exploratory study motivates the development of more advanced models capable of handling non-linear compressible effects beyond our simplified framework.
@@
Overall, this work highlights the Bingham slurry parametric study, and also the advantages of using advanced numerical techniques, such as FEM, to accurately capture the complex interplay of rheological properties, flow dynamics, and operational changes in pipelines transporting non-Newtonian fluids. However, our work also presents some limitations important to highlight and that need to be addressed by future studies on the topic. On the one hand, it is necessary to compare these theoretical predictions with real measurements obtained from physical models—or better yet, with transient measurements from large-scale piping systems. Only in this way can we assess the accuracy of the dynamic model proposed in this article. On the other hand, future research could further refine these models through machine learning approaches for rapid parameter optimization, and stabilization techniques could be explored to enhance model robustness. Concerning the governing equations, future research will also encompass the implementation of a fully unsteady friction model for the pipe walls [16]. This work lays a foundation for more resilient pipeline designs and informed operational strategies, particularly in industries where non-Newtonian fluids pose unique challenges to system stability and durability.
5 Acknowledgments
The authors acknowledge the School of Civil Engineering for providing the facilities to conduct this research. Felipe Galarce acknowledges the financial support from the funding ANID Fondecyt Regular 1250287, DI Vinci Iniciación PUCV 039.731/2025, and the Cyprus Research and Innovation Foundation: Horizon Europe - 2nd Opportunity OPPTY-MSCA/0125. Francisco Martinez acknowledges the financial support from the funding Puntaje de Corte, DI Investigación PUCV 210.703/2024.
6 Appendix: List of designations and symbols used in this document
The following list includes the notation for all variables used in the manuscript:
$x$ streamwise position along the pipeline
$t$ time
$u(x,t)$ mean flow velocity at position $x$, at time $t$
$p(x,t)$ mean flow pressure at position $x$, at time $t$
[…]
$P_{o}$ steady-state pressure scale ($=\rho_{m}gH$)
$c$ shockwave celerity
$g$ acceleration of gravity
$C_{V}$ mixture volume concentration
$\rho_{f}$ fluid density
$\rho_{s}$ solids density
$\rho_{m}$ mixture density
$\gamma$ mixture specific gravity ($=\rho_{m}g$)
$\tau_{y}$ mixture yield stress
$\eta$ Bingham dynamic viscosity
$\Delta P$ Joukowsky overpressure pulse
$H$ reservoir total head
Re Reynolds number
He Hedstrom number
$L$ pipeline length
$D$ pipeline inner diameter
$e$ pipeline wall thickness
$E$ pipeline elasticity
$K_{f}$ fluid compressibility modulus
$K_{s}$ solid compressibility modulus
$K_{m}$ mixture compressibility modulus
$f$ Fanning friction coefficient
$S(u)$ frictional term
$g(t)$ valve closure law
$\Delta x$ spatial step
$\Delta t$ time step
$\Delta a^{*}$ dimensionless version of generic variable $a$
$P_{A}$ maximum allowable pressure of the pipeline
$P_{v}$ vacuum pressure